What is the 50th term of the progression 3, 5, 7, 9


1
Expert's answer
2021-09-14T14:15:15-0400

Let's find the first term and the common difference of this arithmetic progression.

Here, the first term "a=3", the common difference "d=5-3=2."

Then, according to the general form of the arithmetic progression, we get:


"a_n=a+(n-1)d,""a_n=3+(n-1)2,""a_n=2n+1."

Finally, we can find the 50-th term of the arithmetic progression:


"a_n=2\cdot50+1=101."

Answer:

"a_n=101."